THE ETHYL RADICAL IN SUPERFLUID HELIUM NANODROPLETS: ROVIBRATIONAL SPEC-TROSCOPY AND AB INITIO CALCLUATIONS

PAUL L. RASTON, \underline{CHRISTOPHER P. MORADI}, Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556; JAY AGARWAL, JUSTIN. M. TURNEY, HENRY F. SCHAE-FER III, Center for Computational Chemistry, University of Georgia, Athens, Georgia 30602-2556; GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556.

The ethyl radical has been isolated and spectroscopically characterized in ^{4}He nanodroplets. The five fundamental CH stretch bands are observed near 3 μm and have band origins shifted < 1 cm^{−1}from those reported for the gas phase species.^{a}^{,}^{b} The symmetric CH_{2} stretching band (ν_{1}) is rotationally resolved, revealing nuclear spin statistical weights predicted by G_{12} permutation-inversion group theory. A permanent electric dipole moment of 0.28 (2) D is obtained via the Stark spectrum of the ν_{1} band. The four other CH stretch fundamental bands are broadened in helium droplets and lack rotational fine structure. The approximately 1-2 cm^{−1}line widths for these bands are attributed to the homogeneous broadening associated with solvent-mediated rovibrational relaxation dynamics. In addition to these five fundamentals, three A′_{1} overtone/combination bands are observed and have resolved rotational substructure. These are assigned to the 2ν_{12}, ν_{4}+ν_{6}, and 2ν_{6} bands through comparisons to anharmonic frequency computations at the CCSD(T)/cc-pVTZ level of theory.

aS. Davis, D. Uy, D. J. Nesbitt. J. Chem. Phys. 112, 1823-1834 (2000).

bT. Haber, A. C. Blair, D. J. Nesbitt, M. D. Schuder. J. Chem. Phys. 124, 054316 (2006).